Find two consecutive inteqerss whose sum iis 123
Answer by checkley77(12844)

x+x+1=123
2x=123-1
2x=122
x=122/2
x=61 ans. for the lesser integer.
61+1=62 ans. for the larger interger.
Proof:
61+62=123
123=123